Advertisements
Advertisements
Question
If `2/11` is the probability of an event, what is the probability of the event ‘not A’.
Advertisements
Solution
P(A) = `2/11`
P(not A) = P (A’) = 1 – P(A)
= `1 - 2/11`
= `9/11`
APPEARS IN
RELATED QUESTIONS
A letter is chosen at random from the word ‘ASSASSINATION’. Find the probability that letter is
- a vowel
- an consonant
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P(not A).
A and B are events such that P(A) = 0.42, P(B) = 0.48 and P(A and B) = 0.16. Determine P (not B)
In Class XI of a school 40% of the students study Mathematics and 30% study Biology. 10% of the class study both Mathematics and Biology. If a student is selected at random from the class, find the probability that he will be studying Mathematics or Biology.
In an entrance test that is graded on the basis of two examinations, the probability of a randomly chosen student passing the first examination is 0.8 and the probability of passing the second examination is 0.7. The probability of passing at least one of them is 0.95. What is the probability of passing both?
In a certain lottery, 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy one ticket.
A dice is thrown. Find the probability of getting a prime number
A dice is thrown. Find the probability of getting a multiple of 2 or 3.
In a simultaneous throw of a pair of dice, find the probability of getting a doublet
In a simultaneous throw of a pair of dice, find the probability of getting a doublet of prime numbers
In a simultaneous throw of a pair of dice, find the probability of getting an even number on one and a multiple of 3 on the other
In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 6
In a simultaneous throw of a pair of dice, find the probability of getting a sum more than 7
In a simultaneous throw of a pair of dice, find the probability of getting odd number on the first and 6 on the second
In a simultaneous throw of a pair of dice, find the probability of getting a number greater than 4 on each die
In a simultaneous throw of a pair of dice, find the probability of getting a total greater than 8.
In a single throw of three dice, find the probability of getting a total of 17 or 18.
Three coins are tossed together. Find the probability of getting at least two heads
Two dice are thrown. Find the odds in favour of getting the sum 4.
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is white .
A box contains 6 red marbles numbered 1 through 6 and 4 white marbles numbered from 12 through 15. Find the probability that a marble drawn is red or even numbered.
Fill in the blank in the table:
| P (A) | P (B) | P (A ∩ B) | P(A∪ B) |
| \[\frac{1}{3}\] | \[\frac{1}{5}\] | \[\frac{1}{15}\] | ...... |
One of the two events must happen. Given that the chance of one is two-third of the other, find the odds in favour of the other.
A card is drawn at random from a well-shuffled deck of 52 cards. Find the probability of its being a spade or a king.
A card is drawn from a deck of 52 cards. Find the probability of getting an ace or a spade card.
The probability that a leap year will have 53 Fridays or 53 Saturdays is
A and B are two events such that P (A) = 0.25 and P (B) = 0.50. The probability of both happening together is 0.14. The probability of both A and B not happening is
If the probability of A to fail in an examination is \[\frac{1}{5}\] and that of B is \[\frac{3}{10}\] . Then, the probability that either A or B fails is
A bag contains 5 black balls, 4 white balls and 3 red balls. If a ball is selected randomwise, the probability that it is black or red ball is
A box contains 6 nails and 10 nuts. Half of the nails and half of the nuts are rusted. If one item is chosen at random, the probability that it is rusted or is a nail is
Three numbers are chosen from 1 to 20. The probability that they are not consecutive is
In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy two tickets.
In a certain lottery 10,000 tickets are sold and ten equal prizes are awarded. What is the probability of not getting a prize if you buy 10 tickets.
